# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v2

@@ VI-A1 Drifting
Drifting is when a vehicle turns with a high sideslip angle, causing tire saturation and losing direct control of its heading [39]. Stable drifting can be achieved by finding “drift equilibria” with constant yaw rate and sideslip angle of the highly nonlinear dynamics [39, 40, 41, 42]. Drift parking, however, remains challenging due to the highly coupled inputs [43] and reduced controllability as the vehicle specifically must not maintain a drift equilibrium [44].
@@ VI-A2 Reach-Avoid Method Comparison
Our drifting tracking model dynamics (see (122) below) are not control affine, so directly applying a CBF or CLBF control approach is not tractable [45, 46]; future work may apply results for non-control-affine systems [47, 48]. Note, [15] poses a control affine vehicle dynamics model but does not include tire saturation, so cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard HJB value function computation [5, 49]; future work will compare against learning-based HJB [13]. The most similar work is RTD [9, 50], which has not been applied to these extreme dynamics. As a future implementation, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison; then, the only difference with PARC is that RTD implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
@@ VI-B1 Planning and Tracking Models
We use a tracking model based on [51, 40, 42, 39], detailed in Appendix -H. Our tracking controller is based on [51]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
@@ VI-B2 Environment and Demo Setup
The objective of this demo is to drift safely into a parallel parked state between two closely parked cars, as shown in Fig. 1 and Fig. 10. Our environment is inspired by a real-world demo video [52].
@@ VI-B4 Discussion
VII Conclusion
@@
This paper has presented an approach to use Piecewise Affine Reach-avoid Computation (PARC) for safe goal-reaching robot trajectory planning near danger. PARC enables goal-reaching through narrow gaps and in tight obstacle configurations, even with extreme dynamics. The method outperforms a variety of state-of-the-art reach-avoid methods and establishes a new benchmark for safe, extreme vehicle motion planning. Future work will improve PARC’s computation time,

### v3

@@ VI-A1 Drifting
Drifting is when a vehicle turns with a high sideslip angle, causing tire saturation and losing direct control of its heading [goh2016simultaneous]. Stable drifting can be achieved by finding “drift equilibria” with constant yaw rate and sideslip angle of the highly nonlinear dynamics [goh2016simultaneous, goh2020toward, goel2020opening, weber2023modeling]. Drift parking, however, remains challenging due to the highly coupled inputs [leng2023deep] and reduced controllability as the vehicle specifically must not maintain a drift equilibrium [kolter2010probabilistic].
@@ VI-A2 Reach-Avoid Method Comparison
Our drifting tracking model dynamics (see (96) below) are not control affine, so directly applying a CBF or CLBF control approach is not tractable [ames2016control, dawson2023safe]; future work may apply results for non-control-affine systems [son2019safety, seo2022safety]. Note, [dawson2022safe] poses a control affine vehicle dynamics model but does not include tire saturation, so cannot drift. Furthermore, our planning and tracking models are too high-dimensional for standard HJB value function computation [chen2021fastrack, mitchell2008flexible]; future work will compare against learning-based HJB [bansal2021deepreach]. The most similar work is RTD [kousik2020bridging, liu2022refine], which has not been applied to these extreme dynamics. As a future implementation, RTD’s reachable set can use PARC’s formulation to enable a direct drifting comparison; then, the only difference with PARC is that RTD implicitly represents the avoid set. Due to this similarity, we focus on demonstrating PARC’s capabilities as opposed to a head-to-head comparison.
@@ VI-B1 Planning and Tracking Models
We use a tracking model based on [jelavic2017autonomous, goh2020toward, weber2023modeling, goh2016simultaneous], detailed in Appendix -H. Our tracking controller is based on [jelavic2017autonomous]: we first use nonlinear MPC to bring the car into a drifting state, then switch to open-loop control to complete the parking action (since the vehicle loses controllability when sliding sideways to a stop).
@@ VI-B2 Environment and Demo Setup
The objective of this demo is to drift safely into a parallel parked state between two closely parked cars, as shown in Fig. 1 and Fig. 12. Our environment is inspired by a real-world demo video [guinness2014parallel].
@@ VI-B4 Discussion
VII Hardware Demonstration
Thus far, PARC has only been shown to work in simulation. In reality, the tracking model will have model mismatch with the actual hardware due to disturbances and perception errors. As such, we now present preliminary results for PARC’s reach-avoid guarantees applied on mobile robot hardware navigating a narrow gap, as shown in Fig. 13. Specifically, we use a TurtleBot3 Burger Model [grossimplementation] differential drive robot running ROS 2 [macenski2022robot].
VII-A Environment and Demo Setup
[…]
Figure 13: Problem setup of the hardware demo. A snapshot of the robot’s realized motion and the tracked trajectory (gradient line) from one run of the experiment was shown. The robot successfully avoided the obstacles (red) and reached the goal set (green).
The robot is equipped with IMU sensors and wheel encoders, which gives estimation of its planning states $p_{x}$, $p_{y}$, and $\theta$. As future work, these states can be more accurately estimates using an onboard lidar sensor.
VII-B Planning and Tracking Models
[…]
VII-C Results
[…]
VII-D Discussion
[…]
VIII Conclusion
@@
This paper has presented an approach to use Piecewise Affine Reach-avoid Computation (PARC) for safe goal-reaching robot trajectory planning near danger. PARC enables goal-reaching through narrow gaps and in tight obstacle configurations, even with extreme dynamics. The method outperforms a variety of state-of-the-art reach-avoid methods and establishes a new benchmark for safe, extreme vehicle motion planning. Future work will improve PARC’s computation time, consider more types of uncertainty as opposed to lumped tracking error, and extend beyond mobile robots.